OpenStudy (anonymous):

find the missing factor 2x^2+13x+15=(x+5)(_____)

OpenStudy (anonymous):

(2x + 3)

OpenStudy (anonymous):

The discriminant is \[13^{2} + 4.2.15 = 49\] \[x_{1} = (-13 + \sqrt{49} ) / 2.2 = 3/2\] and \[x_{2} = -5\] So the answer is 3/2